OpenStudy (anonymous):

Which is the best strategy to solve this problem? Mary-Ann is making gift baskets for her friends. She has 32 candles and 64 lotion bottles. She splits the candles and lotion bottles equally among 8 baskets. How many items were in each basket?

OpenStudy (anonymous):

A. Draw a diagram. Draw 8 circles to represent baskets. Think 32 ÷ 8 = 4, so draw 4 candles in each circle. Think 64 ÷ 8 = 8, so draw 8 dots in each circle. Count up the objects in a circle to find the number of items in each basket. B. Guess and test. Guess that the number of items in each basket is 9. Multiply 8 × 9. Ask yourself if this is the same as the sum of 32 + 64? If not, revise your guess, and test your answer again. C. Write a number sentence. (32 + 64) × 8 =

OpenStudy (anonymous):

C not sure

OpenStudy (anonymous):

@K12awesomeness

OpenStudy (shaik0124):

a is the answer

OpenStudy (k12awesomeness):

It's A, because, You need to divide and see how many candles fit in each basket and how many lotion bottles fit evenly.. To do that, you must use a diagram

OpenStudy (anonymous):

thanks guys I told you I wasn't sure lol

OpenStudy (k12awesomeness):

lol.. no problem!

OpenStudy (shaik0124):

bcoz there are 32+64=96 (candels n lotions) in total as there are 8 baskets if we put 12 items in a single basket then 12 items*8 baskets=96 items so each basket will have equal no of items option a gives the answer 4 candles and 8 lotions in each basket=12

OpenStudy (anonymous):

thanks